OpenStudy (anonymous):

A university bookstore recently sold a wirebound graph-paper notebook for $1.10, and a college-ruled notebook for $3.16. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $102.38. How many of each type were sold?

OpenStudy (anonymous):

x=college-rule y=wirebound 3.16x+1.10y=102.38 x+y=50 y=50-x 3.16x+1.10(50-x)=102.38 3.16x+55-1.10x=102.38 2.06x=47.38 x=23 college ruled 23+y=50 y=27 wirebound

OpenStudy (anonymous):

basically set up 2 equations and substitute one in the other.

OpenStudy (amistre64):

wirebound graph-paper $1.10 college-ruled notebook $3.16. 50 sold for a total of $102.38. How many of each type were sold? amount @ 3.16 = (102.38 - 50(1.10))/(3.19-1.10)

OpenStudy (amistre64):

adjust for typos

OpenStudy (amistre64):

23 @ 3.16 27 @ 1.10 is what i get; but dbl chk